Calculate the volume occupied by 8.8 g of CO2 at 31.1°C and 1 bar pressure. R = 0.083 bar L K–1 mol–1

Grade:12

2 Answers

Gaurav
9 years ago
Hello Student
pV = nRT, we can replace n by m/M, where,
m = mass of gas
M = Molecular mass in gmol–1
So, pV = (m/M) X RT,
Hence, V = (mRT)/ Mp
Quantites given are :
m = 8.8 g
R = 0.083 bar LK–1 mol–1
T = 31.1°C = 304.1 K
M = 44 g
p = 1 bar
Plugging these values in above equation we get, V = 5.05 L.
Kushagra Madhukar
3 years ago
Dear student,
Please find the attached solution.
 
pV = nRT, we can replace n by m/M, where,
m = mass of gas
M = Molecular mass in gmol–1
So, pV = (m/M) X RT,
Hence, V = (mRT)/ Mp
Quantites given are :
m = 8.8 g
R = 0.083 bar LK–1 mol–1
T = 31.1°C = 304.1 K
M = 44 g
p = 1 bar
Substituting these values in above equation we get, V = 5.05 L
 
Hope it helps.
Thanks and regards,
Kushagra